OpenStudy (anonymous):

if I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost?

OpenStudy (anonymous):

it c right?

OpenStudy (campbell_st):

2a + 6o = 5.60 (1) 6a + 2o = 4.80 (2) solve via the elimination method 3 x (1) - (2) 6a + 18o = 16.80 - 6a + 2o = 4.80 ---------------- 16o = 12 o = $0.75